The weights (lb) of discarded plastic from a sample of households are 62 with a mean of 1,911 lb and the standard deviation is 1,065 lb. Use a 0,05 significance level to test the claim that the mean weight of discarded plastic from the population is greater than 1,800 lb.
1
Expert's answer
2020-05-13T19:03:10-0400

Null hypothesis "H_0:\mu=1800."

Alternative hypothesis "H_a:\mu>1800."

Test statistic: "z=\frac{\bar x-\mu}{\frac{\sigma}{\sqrt{n}}}=\frac{1911-1800}{\frac{1065}{\sqrt{62}}}=0.82."

P-value: "p=P(Z>0.82)=0.206."

Since the P-value is greater than 0.05, fail to reject the null hypothesis.